### v1

# Streaming algorithms for products of probabilities

Abstract
We consider streaming algorithms for approximating a product of input probabilities up to multiplicative error of $1-\epsilon$. It is shown that every randomized streaming algorithm for this problem needs space $\Omega(\log n+\log b-\log\epsilon)-\mathcal{O}(1)$, where $n$ is length of the input stream and $b$ is the bit length of the input numbers. This matches an upper bound from Alur et al. up to a constant multiplicative factor. Moreover, we consider the threshold problem, where it is asked whether the product of the input probabilities is below a given threshold. It is shown that every randomized streaming algorithm for this problem needs space $\Omega(n\cdot b)$.
keywords:
streaming algorithms
category:
1 Introduction
The streaming model is one of the central concepts in data science to deal with huge data sets, where random access is no longer feasible. It is based on the processing of an input stream of data values. Every data value has to be processed immediately when it arrives. For a good introduction to streaming algorithms see [2, 6]. In this paper we consider the problem of computing probabilities in the streaming model. More precisely, consider a stream of probabilities $q_{1},q_{2},\ldots,q_{n}$ where every $q_{i}$ is given as a fraction $r_{i}/s_{i}$ of integers $r_{i},s_{i}$ with $r_{i}\geq 0$ and $1\leq r_{i}\leq s_{i}$. The goal is to compute a good approximation of the product probability $\prod_{1\leq i\leq n}q_{i}$. This problem has been first studied by Alur et al. in an automata theoretic setting [1]. It was shown that a $(1-\epsilon)$-approximation [^1] A $(1-\epsilon)$-approximation of a number $x$ is any number $x^{\prime}$ with $(1-\epsilon)x<x^{\prime}<x/(1-\epsilon)$. of the product $\prod_{1\leq i\leq n}q_{i}$ can be computed by a deterministic streaming algorithm in space $\log_{2}(n^{2}b/\epsilon)=2\log n+\log b-\log\epsilon$. Here, $n$ is the length of the stream and $b$ is the maximal bit length of the numerators $r_{i}$ and denominators $s_{i}$ in the stream of probabilities $q_{i}=r_{i}/s_{i}$. [^2] The space bound stated in [1] is actually larger than $2\log n+\log b-\log\epsilon$ since the authors in [1] also count the space for internal computations; see Section 2.1 for more details on this.
Our first result is a lower bound that matches the upper bound $22\log n+\log b-\log\epsilon+$ 2 up to a constant factor. This is also true for randomized streaming algorithms that produce a $(1-\epsilon)$-approximation of the product $\prod_{1\leq i\leq n}q_{i}$ with high probability. Our lower bound only holds if $\log(1/\epsilon)\leq b<\log(1/(1-\epsilon))\cdot n/2$ holds. This is a reasonable setting. If we consider the approximation parameter $\epsilon$ as a small constant then the requirement becomes $\Omega(1)\leq b\leq\mathcal{O}(n)$. Since the length $n$ of the stream is usually very large, it is reasonable to assume $b\leq\mathcal{O}(n)$. In a realistic scenario one would even require $b\leq\mathcal{O}(\log n)$.
Our second result is a new lower bound for the threshold problem: the input consists again of a stream of probabilities $q_{1},q_{2},\ldots,q_{n}$, where $q_{i}=r_{i}/s_{i}$ and the integers $r_{i}$ and $s_{i}$ need at most $b$ bits. In addition, there is a threshold probability $t=r/s$, where $r$ and $s$ also need at most $b$ bits. The question is whether $\prod_{1\leq i\leq n}q_{i}<t$ holds. It was shown in [1, Theorem A.1] that every deterministic streaming algorithm for the corresponding problem, where $b=\log n$ and the $q_{i}$ and $t$ are not restricted to the interval $[0,1]$, has to store $\Omega(n)$ many bits. It was asked whether this lower bound also holds when all $q_{i}$ and $t$ are probabilities. We prove that this is indeed the case and moreover improve the lower bound from $\Omega(n)$ to $\Omega(n\cdot b)$ if $b\geq\Omega(\log n)$ for randomized streaming algorithms. Note that the threshold problem can be always solved in space $\Omega(n\cdot b)$ by naively computing the product $\prod_{1\leq i\leq n}q_{i}$. On the other hand, if $n$ is very large compared to $b$, then there is a better (deterministic) algorithm that only needs $\mathcal{O}(2^{b})$ bits; see Theorem 4.1.
In the final section we consider the product approximation problem for rational probabilities in the sliding window setting. Here, the goal is to maintain a $(1-\epsilon)$-approximate of the product of the last $m$ probabilities from the input stream. The parameter $m$ is called the window size. We provide a space lower bound of $\Omega(m\cdot(\log b-\log\epsilon-\Theta(1)))$ for randomized sliding window algorithms, where as before $b$ is the bit length of the input probabilities. Currently, we are not able to match this lower bound by an upper bound. In Section 5 we provide two incomparable upper bounds: $2mb$ (which is trivial) and $m\cdot(\log m+\log b-\log\epsilon)$.
2 Preliminaries
For a possibly infinite set $A$ we denote with $A^{*}$ the set of all finite words (or sequences) over the set $A$. We write $A^{\leq n}$ for the subset of words of length at most $n$. With $\log$ we always denote the logarithm to the base two. For a positive rational number $q=r/s$ with $r,s\in\mathbb{N}$ and $\mathsf{gcd}(r,s)=1$ we define the bit size of $q$ as $\|q\|=\max\{\lceil\log r\rceil,\lceil\log s\rceil\}$. In addition, we define $\|0\|=0$. Let =$\mathbb{P}=\{q\in\mathbb{Q}:0\leq q\leq 1\}$ be set of rational probabilities and $\mathbb{P}_{b}=\{q\in\mathbb{P}:\|q\|\leq b\}$ be the set of rational probabilities of bit size at most $b$.
2.1 Streaming algorithms
We only consider streaming problems where the input values are rational probabilities and there is a set of allowed rational output probabilties (this also covers the case, where the output is either 00 or $1$). In our context, the output probabilities will be approximations of some rational probability that is uniquely determined by the input sequence.
Formally, a streaming problem is specified by a function $F:\mathbb{P}^{*}\to 2^{\mathbb{P}}$ mapping a sequence $q_{1}q_{2}\cdots q_{n}$ of rational probabilities to a set $F(q_{1}q_{2}\cdots q_{n})$ of rational probabilities. Below, we slice the set of all input sequences $\mathbb{P}^{*}$ into the sets $\mathbb{P}_{b}^{\leq n}$ consisting of input sequences of length at most $n$, where in addition all input numbers have bit size at most $b$.
A deterministic streaming algorithm can be formalized by a family of deterministic finite automata $\mathcal{D}=(\mathcal{A}_{n,b})_{n,b\in\mathbb{N}}$, where the automaton $\mathcal{A}_{n,b}$ is responsible for inputs from the slice $\mathbb{P}_{b}^{\leq n}$. Hence, the input alphabet of $\mathcal{A}_{n,b}$ is $\mathbb{P}_{b}$. Let $Q_{n,b}$ be the set of states of $\mathcal{A}_{n,b}$. Its elements can be seen as the memory states of the streaming algorithm. There is a distinguished initial state in $\mathcal{A}_{n,b}$. The dynamics of the streaming algorithm is described by the state transition function $\delta_{n,b}:Q_{n,b}\times\mathbb{P}_{b}\to Q_{n,b}$ such that for a memory state $s\in Q_{n,b}$ and a new data value $q\in\mathbb{P}_{b}$, $\delta_{n,b}(s,q)\in Q_{n,b}$ is the new memory state. In addition, $\mathcal{A}_{n,b}$ has an output function $o_{n,b}:Q_{n,b}\to\mathbb{P}$ instead of a set of final states. Such an automaton computes a function $f_{\mathcal{A}_{n,b}}:\mathbb{P}_{b}^{\leq n}\to\mathbb{P}$ as follows: Starting from the initial state, the automaton reads the input stream $w\in\mathbb{P}_{b}^{*}$. After reading the last input number from $w$, the automaton is in a certain state $s\in Q_{n,b}$. The output value $f_{\mathcal{A}_{n,b}}(w)$ is then $o_{n,b}(s)$. A deterministic streaming algorithm $\mathcal{D}=(\mathcal{A}_{n,b})_{n,b\in\mathbb{N}}$ for the streaming problem $F$ has to satisfy $f_{\mathcal{A}_{n,b}}(w)\in F(w)$ for all $n,b\in\mathbb{N}$ and $w\in\mathbb{P}_{b}^{\leq n}$. Finally, the space complexity of the streaming algorithm $\mathcal{D}$ is the mapping $\mathsf{space}_{\mathcal{D}}(n,b)=\lceil\log_{2}|Q_{n,b}|\rceil$, which is the number of bits required to encode the states in binary notation.
Some comments regarding our definition are appropriate. First, our definition of a deterministic streaming algorithm is non-uniform in the sense that for every input slice $\mathbb{P}_{b}^{\leq n}$ we have a separate algorithm (the automaton $\mathcal{A}_{n,b}$). In principle the function $(n,b)\mapsto\mathcal{A}_{n,b}$ may be not computable. In a real streaming algorithm, this function will be certainly computable. On the other hand, the main focus in this paper is on lower bounds, and lower bounds for non-uniform algorithms also hold for uniform algorithms. Related to this point is the fact that we ignore in our definition of the space complexity $\mathsf{space}_{\mathcal{D}}(n,b)=\lceil\log_{2}|Q_{n,b}|\rceil$ the space needed for computing the the mapping $\delta_{n,b}$, i.e., the space needed to compute the next memory state from the previous memory state and the input data value (this space is what we called the space for internal computations in footnote 2). In general, this space will also grow with $n$ and $b$. The justification for ignoring this space is the same as for non-uniformity. A lower bound for our measure of space complexity also holds for the more realistic measure that incorporates the internal space needed for computing the state transformation mapping. Another point is that lower bound techniques for streaming algorithms (which are typically based on information theoretic arguments or communication complexity) only yield lower bounds on our space complexity.
Randomized streaming algorithms can be defined analogously to deterministic streaming algorithms by taking for every $\mathcal{A}_{n,b}$ a probabilistic finite automaton (see for details on probabilistic automata). For an input stream $w\in\mathbb{P}_{b}^{*}$, we then obtain a random variable $\mathsf{R}_{\mathcal{A}_{n,b}}^{w}$, whose support is the set of all output values $o_{n,b}(s)$ for $s\in Q_{s,b}$. For an output value $q$, $\mathsf{R}_{\mathcal{A}_{n,b}}^{w}(q)$ is the probability that on input $w$ the probabilistic automaton $\mathcal{A}_{n,b}$ returns $q$. A randomized streaming algorithm $\mathcal{R}=(\mathcal{A}_{n,b})_{n,b\in\mathbb{N}}$ for the streaming problem $F$ has to satisfy
$\mathsf{Prob}[\mathsf{R}_{\mathcal{A}_{n,b}}^{w}\in F(w)]\geq\frac{2}{3}$
for all $n,b\in\mathbb{N}$ and $w\in\mathbb{P}_{b}^{\leq n}$. The probability $\frac{2}{3}$ can be replaced by any $1-\lambda$ by running $\mathcal{O}(\log(1/\lambda))$ many independent copies of $\mathcal{A}_{n,b}$.
We consider the following streaming problems in this paper, where $10\leq\epsilon<$ 1 is an approximation ratio and $q_{1},q_{2},\ldots,q_{n}\in\mathbb{P}$:
  • The streaming problem $\mathsf{APP}_{\epsilon}$ (approximation of products of probabilities) is
    $\mathsf{APP}_{\epsilon}(q_{1}q_{2}\cdots q_{n})=\bigg{\{}q\in\mathbb{P}:(1-\epsilon)\prod_{i=1}^{n}q_{i}<q<\frac{1}{1-\epsilon}\prod_{i=1}^{n}q_{i}\bigg{\}}.$
  • The threshold problem for products of probabilities, $\mathsf{TPP}$ for short, is
    $\mathsf{TPP}(q_{1}q_{2}\cdots q_{n})=\begin{cases}1&\text{ if }\prod_{i=2}^{n}q_{i}<q_{1}\\ 0&\text{ otherwise. }\end{cases}$
2.2 Communication complexity
Lower bounds for randomized streaming algorithms are typically proved by using lower bounds on randomized communication complexity; see [4, 8] for a detailed introduction. We only use the one-way setting where Alice sends a single message to Bob.
Consider a function $f\colon X\times Y\to\{0,1\}$ for some finite sets $X$ and $Y$, which we call a communication problem. A randomized one-way (communication) protocol $P$ for $f$ consists of two parties called Alice and Bob. The input for Alice (resp., Bob) is an element $x\in X$ and a random choice $r\in R$ (resp., $y\in Y$ and a random choice $s\in S$). Here, $R$ and $S$ are finite sets and the random choices of Alice and Bob are independent. The goal of Alice and Bob is to compute $f(x,y)$. For this, Alice computes from here input $x\in X$ and random choice $r\in R$ a message $a(x,r)\in\{0,1\}^{*}$ which is send to Bob. Bob then computes from $a(x,r)$, his input $y\in Y$ and random choice $s\in S$ the final output $b(a(x,r),y,s)\in\{0,1\}$ of the protocol. We also assume a probability distribution on the set $R$ (resp., $S$) of Alice’s (resp., Bob’s) random choices. The protocol $P$ computes $f$ if for all $(x,y)\in X\times Y$ we have
$\mathsf{Prob}_{r\in R,s\in S}[b(a(x,r),y,s)=f(x,y)]\geq\frac{2}{3}.$ (1)
The cost of the protocol is the maximum length $|a(x,r)|$ taken over all $(x,r)\in X\times R$. The randomized one-way communication complexity of $f$ is the minimal cost among all (one-way) randomized protocols that compute $f$. Here, the size of the finite sets $R$ and $S$ is not restricted. The choice of the constant $2/3$ in (1) is arbitrary in the sense that changing the constant to any $1-\lambda$ only changes the communication complexity by a constant (depending on $\lambda$), see [4, p. 30]. Also note that we only use the private version of randomized communication protocols, where Alice and Bob make private random choices from the sets $R$ and $S$, respectively, and their choices are not known to the other party (in contrast to the public version of randomized communication protocols).
We will use the following communication problems in this paper: The greater-than problem $\mathsf{GT}_{m}$ is defined as follows:
  • Alice’s input is a number $a\in\{1,\dots,m\}$.
  • Bob’s input is a number $b\in\{1,\dots,m\}$.
  • Bob’s goal is to determine whether $a>b$.
The index-greater-than problem $\mathsf{IndexGT}_{m,n}$ is defined as follows:
  • Alice’s input is a sequence $a_{1}a_{2}\cdots a_{n}$ with $a_{i}\in\{1,\dots,m\}$ for all $i$.
  • Bob’s input is a number $b\in\{1,\dots,m\}$ and a number $i\in\{1,\ldots,n\}$.
  • Bob’s goal is to determine whether $a_{i}<b$.
Theorem 2.1.
The following hold:
  • The randomized one-way communication complexity of $\mathsf{GT}_{m}$ is $\Theta(\log m)$ [5].
  • The randomized one-way communication complexity of $\mathsf{IndexGT}_{m,n}$ is $\Theta(n\cdot\log m)$ [3].
3 Lower bound for the product approximation problem
In this section we consider the problem $\mathsf{APP}_{\epsilon}$. The following upper bound was shown in [1].
Theorem 3.1 ([1]).
For every $20<\epsilon<1/$ 2 there is a deterministic streaming algorithm for $\mathsf{APP}_{\epsilon}$ with space complexity $\log(n^{2}b/\epsilon)=2\log n+\log b-\log\epsilon$.
Let us briefly explain the idea since it will be needed later. Assume that the input stream is $q_{1}q_{2}\cdots q_{n}$. Define $\epsilon^{\prime}=\epsilon/n$. For an $a\in\mathbb{N}$ we define the bucket $B_{a,\epsilon^{\prime}}=((1-\epsilon^{\prime})^{a+1},(1-\epsilon^{\prime})^{a}]$. Then, the algorithm computes for every input number $q_{i}$ the unique $a_{i}\in\mathbb{N}$ such that $q_{i}\in B_{a_{i},\epsilon^{\prime}}$. Then the streaming algorithmen for $\mathsf{APP}_{\epsilon}$ computes the sum $a=\sum_{i=1}^{n}a_{i}$ and at the end returns $(1-\epsilon^{\prime})^{a}$. It can be shown that this is a $(1-\epsilon)$-approximation of the product of the $q_{i}$. Moreover, every bucket index $a_{i}$ can be bounded by $nb/\epsilon$ (in [1] the bound $4nb/\epsilon$ is stated, but it easy to see that $nb/\epsilon$ suffices.) Hence, every partial sum of the $a_{i}$ is bounded by $n^{2}b/\epsilon$ for which $2\log n+\log b-\log\epsilon$ bits suffice.
The space bound stated in [1] also contains the space needed for computing the memory state transition function (called $\delta_{n,b}$ in Section 2.1). This is mainly the space for computing the bucket indices $a_{i}$. Recall from Section 2.1 that we ignore this internal space in our model. We match the upper bound $2\log n+\log b-\log\epsilon$ by a lower bound up to a factor of two:
Theorem 3.2.
Let $20<\epsilon<1/$ 2. Every deterministic streaming algorithm for $\mathsf{APP}_{\epsilon}$ has space complexity at least $\log n+\log b-\log\epsilon-\Theta(1)$ whenever $-\log\epsilon\leq b<-\log(1-\epsilon)\cdot n/2$.
Proof 3.3.
Fix an $\epsilon$ with $20<\epsilon<1/$ 2 and let $\delta=-\log(1-\epsilon)$. Note that $10<\delta<$ 1. Let $\mathcal{D}$ be a deterministic streaming algorithm for $\mathsf{APP}_{\epsilon}$. It suffices to show that $\mathsf{space}_{\mathcal{D}}(2n,b)\geq\log n+\log b-\log\delta-\Theta(1)$ under the assumption
$-\log\epsilon\leq b<\delta\cdot n.$ (2)
Clearly, we can then subsitute $n$ by $n/2$. Moreover, it can be easily shown that the difference $\log\epsilon-\log\delta=\log\epsilon-\log\log(1/(1-\epsilon))$ belongs to the interval $(-1,\log\ln 2)\subseteq(-1,-0.52)$ when $20<\epsilon<1/$ 2. Thus, we have $\log\epsilon-\Theta(1)=\log\delta-\Theta(1)$.
Due to (2) (which yields 1/$21/2^{b}\leq\epsilon<1/$ 2) there is an integer $k$ with $3\leq k\leq 2^{b}$ and
$\frac{1}{k}\leq\epsilon\leq\frac{1}{k-1}$ (3)
and hence
$\frac{k-2}{k-1}\leq 1-\epsilon\leq\frac{k-1}{k}.$ (4)
We consider streams of length at most $2n$ of rational numbers from the set
$S=\bigg{\{}\frac{k-1}{k},\frac{k-2}{k-1},2^{-b}\bigg{\}}.$ (5)
Note that $\|q\|\leq b$ for all $q\in S$. Let us define
$y:=\left\lfloor\frac{nb}{\delta}\right\rfloor.$
For $j\geq 0$ we define the $j$-th bucket as the interval
$B_{j}=\big{(}(1-\epsilon)^{j+1},(1-\epsilon)^{j}\big{]}\subseteq(0,1].$
Note that every real number $r\in(0,1]$ belongs to a unique bucket.
Claim 1.
For every $j$ with $0\leq j\leq y$ there are numbers $q_{1},q_{2},\ldots,q_{\ell}\in S$ with $\ell\leq 2n$ and $\prod_{i=1}^{\ell}q_{i}\in B_{j}$.
{claimproof}
Let $0\leq j\leq y$. Choose the number $m\geq 0$ maximal such that $2^{-bm}\geq(1-\epsilon)^{j}$, which is equivalent to $2^{bm}\leq 2^{\delta j}$, i.e., $bm\leq\delta j$. We therefore have
$m=\left\lfloor\frac{j\delta}{b}\right\rfloor\leq\frac{j\delta}{b}\leq\frac{y\delta}{b}\leq n.$
We also have
[…]
$=$ $(1-\epsilon)^{-n}(1-\epsilon)^{j}\ =\ (1-\epsilon)^{j-n}.$
It follows that if $B_{s}$ is the unique bucket containing $2^{-bm}$ then $\max\{0,j-n\}\leq s\leq j$. Hence, we have $0\leq j-s\leq n$. We now define the first $m\leq n$ numbers of the stream as $q_{1}=q_{2}=\cdots=q_{m}=2^{-b}$. Hence we have
$q_{1}q_{2}\cdots q_{m}=2^{-bm}\in B_{s},$ (6)
where $s$ satisfies $0\leq j-s\leq n$.
In the second step we choose the numbers $q_{m+1},\ldots,q_{m+j-s}$ inductively such that $q_{1}q_{2}\cdots q_{m+j-s}\in B_{j}$. Note that $m+j-s\leq 2n$. Assume that $q_{m+1},\ldots,q_{m+d}\in S$ have been chosen for some $0\leq d<j-s$ such that $q_{1}q_{2}\cdots\cdots q_{m+d}\in B_{s+d}$. Note that for $d=0$ this holds by (6). We choose $q_{m+d+1}$ by a case distinction depending on $q_{1}q_{2}\cdots\cdots q_{m+d}$. It might actually happen that $q_{1}q_{2}\cdots\cdots q_{m+d}$ satisfies both cases 1 and 2 below. Then we can choose $q_{m+d+1}$ according to one of them. Recall also that $k$ is such that $3\leq k\leq 2^{b}$ and (4) holds.
Case 1. We have
$q_{1}q_{2}\cdots q_{m+d}\in\left((1-\epsilon)^{s+d+1},\frac{k}{k-1}(1-\epsilon)^{s+d+1}\right]\subseteq B_{s+d}.$
Inclusion in $B_{s+d}$ holds since $k(1-\epsilon)/(k-1)\leq 1$ by (4). Then we get
$q_{1}q_{2}\cdots q_{m+d}\cdot\frac{k-1}{k}$$\in$ $\left(\frac{k-1}{k}(1-\epsilon)^{s+d+1},(1-\epsilon)^{s+d+1}\right]$
$\subseteq$ $\left((1-\epsilon)^{s+d+2},(1-\epsilon)^{s+d+1}\right]=B_{s+d+1}$
where the inclusion follows from $(k-1)/k\geq(1-\epsilon)$; see (4). We can therefore choose $q_{m+d+1}=(k-1)/k$.
Case 2. We have
$q_{1}q_{2}\cdots q_{m+d}\in\left(\frac{k-1}{k-2}(1-\epsilon)^{s+d+2},(1-\epsilon)^{s+d}\right]\subseteq B_{s+d},$
where the inclusion in $B_{s+d}$ follows from $(k-1)(1-\epsilon)/(k-2)\geq 1$; see again (4). Also note that $k\geq 3$. Then we get
$q_{1}q_{2}\cdots q_{m+d}\cdot\frac{k-2}{k-1}$$\in$ $\left((1-\epsilon)^{s+d+2},\frac{k-2}{k-1}(1-\epsilon)^{s+d}\right]$
$\subseteq$ $\left((1-\epsilon)^{s+d+2},(1-\epsilon)^{s+d+1}\right]=B_{s+d+1},$
where the inclusion follows again from (4). We can therefore choose $q_{m+d+1}=(k-2)/(k-1)$.
It remains to show that the cases 1 and 2 cover all possibilities, i.e., that
$\frac{k}{k-1}(1-\epsilon)^{s+d+1}\geq\frac{k-1}{k-2}(1-\epsilon)^{s+d+2}.$
After cancelling $(1-\epsilon)^{s+d+1}$ we obtain
$\frac{k}{k-1}\geq\frac{k-1}{k-2}(1-\epsilon)=\frac{k-1}{k-2}-\epsilon\cdot\frac{k-1}{k-2}\,,$
or, equivalently,
$\epsilon\geq\frac{k-2}{k-1}\cdot\bigg{(}\frac{k-1}{k-2}-\frac{k}{k-1}\bigg{)}=\frac{k-2}{k-1}\cdot\frac{1}{(k-1)(k-2)}=\frac{1}{(k-1)^{2}}\ .$
But this is true, because by (3) and $k\geq 3$ we have $\epsilon\geq k^{-1}\geq(k-1)^{-2}$. This concludes the proof of Claim 1. Consider the buckets $B_{3j}$ for $0\leq j\leq\lfloor y/3\rfloor$. By Claim 1, for every $0\leq j\leq\lfloor y/3\rfloor$ there is a number $a_{j}\in B_{3j}$ such that $a_{j}$ can be written as a product of length at most $2n$ over the set $S$. Let $s_{j}$ be the corresponding input stream whose product is $a_{j}$.
Consider now $i$ and $j$ with $0\leq i<j\leq\lfloor y/3\rfloor$. We have $a_{j}<a_{i}$. Moreover,
$\frac{a_{j}}{1-\epsilon}\leq\frac{(1-\epsilon)^{3j}}{1-\epsilon}=(1-\epsilon)^{3j-1}\leq(1-\epsilon)^{3i+2}=(1-\epsilon)(1-\epsilon)^{3i+1}<(1-\epsilon)a_{i}.$
This means that the deterministic streaming algorithm $\mathcal{D}$ for $\mathsf{APP}_{\epsilon}$ must yield different output numbers for the input streams $s_{i}$ and $s_{j}$. In particular, the streaming algorithm must arrive in different memory states for the input streams $s_{i}$ and $s_{j}$. Since this holds for all $0\leq i<j\leq\lfloor y/3\rfloor$, the algorithm must have at least $\lfloor y/3\rfloor+1\geq y/3$ memory states and therefore must store at least $\log(y/3)=\log n+\log b-\log\delta-\Theta(1)$ bits.
We now extend Theorem 3.2 to randomized streaming algorithms.
Theorem 3.4.
Let $20<\epsilon<1/$ 2. Every randomized streaming algorithm for $\mathsf{APP}_{\epsilon}$ has space complexity $\Omega(\log n+\log b-\log\epsilon-\Theta(1))$ whenever $-\log\epsilon\leq b<-\log(1-\epsilon)\cdot n/2$.
Proof 3.5.
We can show this result by a standard reduction to the one-way communication complexity of $\mathsf{GT}_{m}$; see Theorem 2.1. Let $\mathcal{R}$ be a randomized streaming algorithm for $\mathsf{APP}_{\epsilon}$. As in the proof of Theorem 3.2, we define $\delta=-\log(1-\epsilon)$. It suffices to show that $\mathsf{space}_{\mathcal{R}}(2n,b)\geq\Omega(\log n+\log b-\log\delta-\Theta(1))$ under the assumption (2). We can assume that the error probability of $\mathcal{R}$ is bounded by $1/5$. Consider the number $y$ and the buckets $B_{5i}$ for $0\leq i\leq\lfloor y/5\rfloor$ from the proof of Theorem 3.2. Let $s_{i}$ be an input stream whose product yields a number $a_{i}\in B_{5i}$.
We show that $\mathcal{R}$ yields a randomized one-way protocol for $\mathsf{GT}_{\lfloor y/5\rfloor}$. Assume that $i$ is the input for Alice and $j$ is the input for Bob, where $0\leq i,j\leq\lfloor y/5\rfloor$. Alice starts by running the streaming algorithm $\mathcal{R}$ on the input stream $s_{i}$. Thereby she uses here random bits to simulate the probabilistic choices of $\mathcal{R}$. Let $\alpha_{i}$ by the final memory state of $\mathcal{R}$ obtained by Alice. She sends $\alpha_{i}$ to Bob. Bob then runs $\mathcal{R}$ on input $s_{j}$; let $\alpha_{j}$ be the final memory state of $\mathcal{R}$ obtained by Bob. Bob then applies the output function of $\mathcal{R}$ to the memory states $\alpha_{i}$ and $\alpha_{j}$ and obtains rational numbers $r_{i}$ and $r_{j}$ respectively. With probability at least $(5/6)^{2}\geq 2/3$ the output numbers $r_{i}$ and $r_{j}$ satisfy $a_{i}(1-\epsilon)\leq r_{i}\leq a_{i}/(1-\epsilon)$ and $a_{j}(1-\epsilon)\leq r_{j}\leq a_{j}/(1-\epsilon)$, and thus $r_{i}\in B_{5i+1}\cup B_{5i}\cup B_{5i-1}$ and $r_{j}\in B_{5j+1}\cup B_{5j}\cup B_{5j-1}$. This allows Bob to distinguish the cases $i<j$ and $i\geq j$:
  • If $i<j$ then $r_{i}>r_{j}$ and between the unique buckets containing $r_{j}$ and $r_{i}$ there are at least two buckets, namely $B_{5i+2},\ldots,B_{5j-2}$.
  • If $i>j$ then $r_{i}<r_{j}$ and between the unique buckets containing $r_{i}$ and $r_{j}$ there are at least two buckets, namely $B_{5j+2},\ldots,B_{5i-2}$.
  • If $i=j$ then $r_{i},r_{j}\in B_{5i+1}\cup B_{5i}\cup B_{5i-1}$ and there is at most one bucket between the unique buckets containing $r_{i}$ and $r_{j}$.
This shows that the communicated memory state $\alpha_{i}$ must have bit length $\Omega(\log(y/5))=\Omega(\log n+\log b-\log\delta-\Theta(1))$.
4 The threshold problem
In this section we consider the threshold problem $\mathsf{TPP}$. We start with two simple upper bounds:
Theorem 4.1.
Let $B=2^{b}$. There are deterministic streaming algorithms for $\mathsf{TPP}$ with space complexity $2nb$, respectively $\mathcal{O}(B)$.
Proof 4.2.
For notational convenience we consider streams of length $n+1$. Consider such an input stream $q_{0}q_{1}\cdots q_{n}\in\mathbb{P}^{*}$ of $n+1$ rational probabilities with $\|q_{i}\|\leq b$ for all $0\leq i\leq n$. Recall that $q_{0}$ is the threshold value. A streaming algorithm can simply store all the $q_{i}$ with at most $2(n+1)b$ bits in total.
Another way to store the product $\prod_{i=1}^{k}q_{i}$ for $1\leq k\leq n$ is to store for every prime $p\leq B$ how often it appears in the product, where an occurrence in the denominator is counted negative. There are $\mathcal{O}(B/\log B)=\mathcal{O}(B/b)$ primes $p\leq B$ and the total number of occurrences of a prime $p$ in the product is between $-nb$ and $nb$. This leads to the space bound $\mathcal{O}(B(\log n+\log b)/b)$.
This can be improved to $\mathcal{O}(B)$ as follows: We can assume that $q_{0}>0$, otherwise one only has to check whether one of the $q_{i}$ ($1\leq i\leq n$) is zero. Since $\|q_{0}\|\leq b$ we must have $q_{0}\geq 1/B$. Moreover, we can ignore all $1$’s in the stream $q_{1}q_{2}\cdots q_{n}$. If $q_{i}<1$ then $q_{i}\leq(B-1)/B$. Assume now that the number of $i$ with $1\leq i\leq n$ and $q_{i}\leq(B-1)/B$ is at least $B\ln B$. Then, the product $\prod_{i=1}^{n}q_{i}$ is upper bounded by
$\left(\frac{B-1}{B}\right)^{B\ln B}<\left(\frac{1}{e}\right)^{\ln B}=\frac{1}{B}\leq q_{0}.$
Hence, the algorithm can safely output $1$.
This means that we can replace in the above space bound $\mathcal{O}(B(\log n+\log b)/b)$ the value $n$ by $B\ln B\leq\mathcal{O}(Bb)$. We obtain the space bound $\mathcal{O}(B)$ (this also includes the space for storing $q_{0}$).
Theorem 4.3.
Fix a (small) constant $\gamma>0$. Every deterministic streaming algorithm $\mathcal{D}$ for $\mathsf{TPP}$ satisfies $\mathsf{space}_{\mathcal{D}}(2n,(2+\gamma)(\log_{2}n+b))\geq nb$ for $n$ and $b$ large enough.
Proof 4.4.
Let $B=2^{b}$. We take the first $nB+n+1$ prime numbers $p_{1},p_{2},...,p_{nB+n+1}$. It is known that $p_{k}\leq k\cdot(\ln k+\ln\ln k)$ for $k\geq 6$ [7, 3.13]. In all numbers appearing in the two input streams that we will construct, the numerator and denominator are bounded by $p_{nB+n+1}^{2}$. Therefore, the bit size of all the numbers involved is bounded by
$\lceil\log_{2}p_{nB+n+1}^{2}\rceil=\lceil 2\log_{2}p_{nB+n+1}\rceil\leq(2+\gamma)(\log_{2}n+b)$
if $n$ and $b$ are large enough (depending on the constant $\gamma$).
For every $1\leq i\leq n$ we define the set
$\begin{matrix}Q_{i}=\big{\{}\frac{p_{j}}{p_{j+1}}:(i-1)(B+1)+1\leq j\leq i(B+1)-1\big{\}}.\end{matrix}$
Thus, $Q_{i}$ consists of $B$ fractions of two consecutive primes. We then consider the set of all words
$S_{n,b}=\{q_{1}q_{2}\cdots q_{n}\in\mathbb{P}^{n}:q_{i}\in Q_{i}\text{ for all}1\leq i\leq n\}.$
Clearly, $|S_{n,b}|=B^{n}$. We set the threshold (the first number in the stream) to
$t=\frac{p_{1}}{p_{nB+n+1}}.$
Assume now that there is a deterministic streaming algorithm $\mathcal{D}$ for $\mathsf{TPP}$ such that $\mathsf{space}_{\mathcal{D}}(2n,(2+\gamma)(\log_{2}n+b))<n\cdot b=\log_{2}|S_{n,b}|$ bits for $n$ and $b$ large enough. Then there must exist two different words $t\,q_{1}q_{2}\cdots q_{n},\,t\,r_{1}r_{2}\cdots r_{n}\in S_{n,b}$ such that after reading these two words, $\mathcal{D}$ must be in the same memory state. W.l.o.g. we assume that $\prod_{i=1}^{n}q_{i}<\prod_{i=1}^{n}r_{i}$.
We continue the words $tq_{1}q_{2}\cdots q_{n}$ and $tr_{1}r_{2}\cdots r_{n}$, respectively, with the length-$n$ word $(s_{1}/r_{1})(s_{2}/r_{2})\cdots(s_{n}/r_{n})$, where
$s_{i}=\frac{p_{(i-1)(B+1)+1}}{p_{i(B+1)+1}}.$
Note that $s_{i}/r_{i}\leq 1$ and that
$\prod_{i=1}^{n}s_{i}=\prod_{i=1}^{n}\frac{p_{(i-1)(B+1)+1}}{p_{i(B+1)+1}}=t.$
Finally, we have
$\prod_{i=1}^{n}q_{i}\prod_{i=1}^{n}r_{i}^{-1}\prod_{i=1}^{n}s_{i}$ $<$ $\prod_{i=1}^{n}s_{i}=t\text{ and }$ (7)
$\prod_{i=1}^{n}r_{i}\prod_{i=1}^{n}r_{i}^{-1}\prod_{i}^{n}s_{i}$ $=$ $\prod_{i=1}^{n}s_{i}=t.$ (8)
But this is a contradiction, since $\mathcal{D}$ arrives in the same memory state after reading the two words.
Theorem 4.5.
Fix a constant $\gamma>0$. Every randomized streaming algorithm $\mathcal{R}$ for $\mathsf{TPP}$ satisfies $\mathsf{space}_{\mathcal{R}}(2n,(2+\gamma)(\log_{2}n+b))\geq\Omega(n\cdot b)$.
Proof 4.6.
We transform a randomized streaming algorithm $\mathcal{R}$ for $\mathsf{TPP}$ into a randomized one-way protocol for $\mathsf{GT}_{B^{n}}$.
Take the set of words $S_{n,b}$, the numbers $s_{i}$ and the threshold $t$ from the proof of Theorem 4.3. Let $f:\{1,2,\ldots,B^{n}\}\to S_{n,b}$ be the unique bijection such that for all $1\leq i,j\leq B^{n}$ we have: $i<j$ if and only if the product of the numbers in the word $f(i)$ is strictly smaller than the product of the numbers in the word $f(j)$.
On input $i$, Alice computes the word $f(i)=q_{1}q_{2}\cdots q_{n}$, reads $t\,q_{1}\cdots q_{n}$ into $\mathcal{R}$ and sends the resulting memory state $\alpha$ of $\mathcal{R}$ to Bob. Bob then computes from his input $j$ the word $f(j)=r_{1}r_{2}\cdots r_{n}$ and reads, starting from memory state $\alpha$, the word $(s_{1}/r_{1})(s_{2}/r_{2})\cdots(s_{n}/r_{n})$ into the algorithm. From the final output of $\mathcal{R}$, Bob can decide with probability at least $2/3$ whether $i<j$:
  • If $\prod_{i=1}^{n}q_{i}<\prod_{i=1}^{n}r_{i}$ (i.e., $i<j$) then $\mathcal{R}$ outputs $1$ with probability at least $2/3$ because
    $\prod_{i=1}^{n}q_{i}\prod_{i=1}^{n}\frac{s_{i}}{r_{i}}<\prod_{i=1}^{n}s_{i}=t.$
  • If $\prod_{i=1}^{n}q_{i}\geq\prod_{i=1}^{n}r_{i}$ (i.e., $i\geq j$) then $\mathcal{R}$ outputs 00 with probability at least $2/3$ because
    $\prod_{i=1}^{n}q_{i}\prod_{i=1}^{n}\frac{s_{i}}{r_{i}}\geq\prod_{i=1}^{n}s_{i}=t.$
It follows that the communicated memory state $\alpha$ of $\mathcal{R}$ must have bit length $\Omega(\log B^{n})=\Omega(n\cdot b)$.
Corollary 4.7.
Every randomized streaming algorithm $\mathcal{R}$ for $\mathsf{TPP}$ satisfies $\mathsf{space}_{\mathcal{R}}(n,b)\geq\Omega(n\cdot b)$ whenever $b\geq\Omega(\log_{2}n)$.
The condition $b\geq\Omega(\log_{2}n)$ cannot be completely avoided in Corollary 4.7, since this would contradict the upper bound $\mathcal{O}(2^{b})$ in Theorem 4.1.
Also note that the lower bound for $\mathsf{APP}_{\epsilon}$ in Theorem 3.2 holds for input streams over a set of only 3 numbers (namely the set $S$ in (5)). The lower bound from Corollary 4.7 does not hold for a constant number of probabilities, since we can then store the product with $\mathcal{O}(\log n)$ bits.
5 Sliding window products
In this section we consider the following sliding window variant $\mathsf{SWAPP}_{m,\epsilon}$ of $\mathsf{APP}_{\epsilon}$, where $m$ is the size of the sliding window.
$\mathsf{SWAPP}_{m,\epsilon}(q_{1}q_{2}\cdots q_{n})=\bigg{\{}q\in\mathbb{P}:(1-\epsilon)\prod_{i=n-m+1}^{n}q_{i}<q<\frac{1}{1-\epsilon}\prod_{i=n-m+1}^{n}q_{i}\bigg{\}}$
(here, we set $q_{i}=1$ for $i\leq 0$). A randomized sliding window algorithm for $\mathsf{SWAPP}_{m,\epsilon}$ is a collection of probabilistic finite automata $\mathcal{R}=(\mathcal{A}_{b})_{b\in\mathbb{N}}$ such that for every $b\in\mathbb{N}$ and every input stream $w\in\mathbb{P}_{b}^{*}$ we have:
$\mathsf{Prob}[\mathsf{R}_{\mathcal{A}_{b}}^{w}\in\mathsf{SWAPP}_{m,\epsilon}(w)]\geq\frac{2}{3}.$
As usual for sliding window algorithms, we measure the space complexity of a sliding window algorithm for $\mathsf{SWAPP}_{m,\epsilon}$ in the window size $m$ (instead of the total stream length $n$), the bit size $b$ and the approximation ratio $\epsilon$.
Theorem 5.1.
For $20<\epsilon<1/$ 2, there are deterministic sliding window algorithms for $\mathsf{SWAPP}_{m,\epsilon}$ with space complexity $m\cdot(\log m+\log b-\log\epsilon)$, respectively $2mb$.
Proof 5.2.
The upper bound $2mb$ is clear, since the whole window content can be stored with $2mb$ bits. For the upper bound $m\cdot(\log m+\log b-\log\epsilon)$ we use the algorithm for $\mathsf{APP}_{\epsilon}$ from [1]. Recall the sketch after Theorem 3.1. Our sliding window algorithm stores for a window content $q_{1}q_{2}\cdots q_{m}$ the sequence of bucket indices $a_{1},a_{2},\ldots,a_{m}$, where $q_{i}\in B_{a_{i},\epsilon^{\prime}}$ for $\epsilon^{\prime}=\epsilon/m$. By the analysis in [1], $(1-\epsilon^{\prime})^{a}$ (with $a=\sum_{i=1}^{m}a_{i})$ is a $(1-\epsilon)$-approximation of $\prod_{i=1}^{m}q_{i}$. Moreover, every $a_{i}$ is bounded by $mb/\epsilon$, which yields the space bound from the theorem.
Theorem 5.3.
Fix an $\epsilon\in(0,1/2)$. Every randomized sliding window algorithm for $\mathsf{SWAPP}_{m,\epsilon}$ requires space $\Omega(m\cdot(\log b-\log\epsilon-\Theta(1)))$.
Proof 5.4.
Let $\delta:=-\log(1-\epsilon)\in(0,1)$. Since $\log\epsilon-\Theta(1)=\log\delta-\Theta(1)$ (see the beginning of the proof of Theorem 3.2), it suffices to show the space lower bound $\Omega(m\cdot(\log b-\log\delta-\Theta(1)))$.
Define the following additional integers:
$\alpha=\lceil 4\delta\rceil,\qquad c=\left\lfloor\frac{b}{\alpha}\right\rfloor.$ (9)
Consider the set $A=\{2^{-i\cdot\alpha}:0\leq i\leq c-1\}\subseteq\mathbb{P}$ of size $c$. Also note that $\|q\|\leq c\cdot\alpha=b$ for all $q\in A$. We prove the theorem by a reduction to the communication problem $\mathsf{IndexGT}_{m,n}$.
Assume that we have a randomized sliding window algorithm $\mathcal{R}$ for $\mathsf{SWAPP}_{\epsilon}$. We can assume that the error probability of $\mathcal{R}$ is bounded by $1/6$. We obtain a randomized one-way communication protocol for $\mathsf{IndexGT}_{c,m}$ as follows: The input of Alice is a sequence $a_{1},a_{2},\ldots,a_{m}$ with $a_{i}\in A$ and the input of Bob is an index $1\leq i\leq m$ and a number $a\in A$. He wants to find out whether $a_{i}<a$. Alice sends the memory state of the randomized sliding window algorithm after reading $a_{1}a_{2}\cdots a_{m}$ to Bob. This allows Bob to compute with high probability $(1-\epsilon)$-approximations $P^{\prime}_{1}$ and $P^{\prime}_{2}$ of the products
$P_{1}=a_{i}a_{i+1}\cdots a_{m}\quad\text{ and }\quad P_{2}=a_{i+1}\cdots a_{m}a.$
For this, Bob continues the sliding window algorithm from the memory state obtained from Alice with the input numbers $1,1,\ldots,1$ and $1,1,\ldots,1,a$, respectively, where the number of $1$’s is $i-1$. With probability at least $1-1/3$ the computed approximations $P^{\prime}_{1}$ and $P^{\prime}_{2}$ satisfy
$P^{\prime}_{1}\in((1-\epsilon)P_{1},P_{1}/(1-\epsilon))\quad\text{ and }\quad P^{\prime}_{1}\in((1-\epsilon)P_{2},P_{2}/(1-\epsilon)).$ (10)
Assume for the following that (10) holds. We claim that from the quotient $P^{\prime}_{1}/P^{\prime}_{2}$, Bob can determine whether $a_{i}<a$ holds. Note that $P_{1}/P_{2}=a_{i}/a$. Moreover, (10) implies
$\frac{P^{\prime}_{1}}{P^{\prime}_{2}}\in\bigg{(}(1-\epsilon)^{2}\cdot\frac{a_{i}}{a},\;\frac{1}{(1-\epsilon)^{2}}\cdot\frac{a_{i}}{a}\bigg{)}.$
If $a_{i}\leq a$ then $P^{\prime}_{1}/P^{\prime}_{2}<1/(1-\epsilon)^{2}$. On the other hand, if $a_{i}>a$ then by the choice of the set $B$ we must have $a_{i}/a\geq 2^{\alpha}\geq 2^{4\delta}=1/(1-\epsilon)^{4}$, which yields $P^{\prime}_{1}/P^{\prime}_{2}>1/(1-\epsilon)^{2}$. Hence, from the quotient $P^{\prime}_{1}/P^{\prime}_{2}$, Bob can indeed distinguish the cases $a_{i}>a$ and $a_{i}\leq a$.
By Theorem 2.1, every randomized sliding window algorithm for $\mathsf{SWAPP}_{m,\epsilon}$ must use space $\Omega(m\cdot\log c)=\Omega(m\cdot(\log b-\log\delta-\Theta(1)))$.
6 Open problem
For the space complexity of $\mathsf{SWAPP}_{m,\epsilon}$, there is a gap between the upper bounds in Theorem 5.1 and the lower bound from Theorem 5.3 that we would like to close. For $\mathsf{APP}_{\epsilon}$, a smaller gap (by a multiplicative factor of 2) arises from the upper bound in Theorem 3.1 and the lower bound in Theorem 3.2. One may also ask, whether the condition $-\log\epsilon\leq b<-\log(1-\epsilon)\cdot n/2$ in Theorem 3.2 can be relaxed.